OpenStudy (anonymous):

Simplify, show steps please: (cosx/1+sinx)+(1+sinx/cosx)

OpenStudy (*insert name here*):

Get common denominators

OpenStudy (anonymous):

\[\frac{\cos x+1+\sin x}{ 1+\sin x\cos x } ?\]

OpenStudy (*insert name here*):

Not exactly..

OpenStudy (anonymous):

Im just stuck on this whole equation.

OpenStudy (*insert name here*):

Multiply everything through with a (1+sinx) and cosx

OpenStudy (*insert name here*):

So.. \[\LARGE \frac{cosx}{(1+sinx)}*\frac{cosx}{cosx}+\frac{1+sinx}{cosx}*\frac{1+sinx}{1+sinx}\] \[\LARGE = \frac{cos^2x}{(cosx)(1+sinx)}+\frac{(1+sinx)^2}{(cosx)(1+sinx)}\] Can you finish it?

OpenStudy (anonymous):

I got \[\frac{ 1 }{ \cos x+\cos x \sin x }\]

OpenStudy (*insert name here*):

Did you distribute that (1+sinx)^2?

OpenStudy (anonymous):

ok i had my dad do it and his answer was. cscx